OpenStudy (anonymous):

solve the given logarithmic or exponential equation. 1. log_27 x=-2/3 2. log_5 (x+9)-log_5 x=3 can someone include the work? i want to know how to solve these question by myself :)

hartnn (hartnn):

for 1. use the definition of log \(\huge \log_ab=c \implies b = a^c \)

hartnn (hartnn):

for 2. , use this property first \(\large \log_ab-\log_ac =\log_a (b/c)\) for the left part.

OpenStudy (anonymous):

for i got no answer. 1. 1=-2/3log(27) 2. 1≠2/3log(27) 3. no solution is that right?

hartnn (hartnn):

\(\huge \log_ab=c \implies b = a^c \\ \huge \log_{27}x= -2/3 \implies x= 27^{-2/3}\) can you simplify the right side ?

OpenStudy (anonymous):

x=1/9

hartnn (hartnn):

thats correct! :)

hartnn (hartnn):

for the left side of 2nd., use this property \(\large \log_ab-\log_ac =\log_a (b/c)\)

OpenStudy (anonymous):

how do i set that up?

hartnn (hartnn):

\(\large \log_ab-\log_ac =\log_a (b/c) \\ \large \log_5 (x+9)- \log_5 x = ... ?\)

OpenStudy (anonymous):

oh ok. i got 5^(x-1)

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

where does x-1 come from ?

OpenStudy (anonymous):

oh is it just 5?

OpenStudy (anonymous):

wait... should the answer be 9?

hartnn (hartnn):

 see if you can understand that